A bullet leaves the muzzle of a hunter's rifle with the velocity of 460ms-1, 2.75x10-3s after the rifle is fired. What is the length of the muzzle?
1
Expert's answer
2018-03-20T11:01:07-0400
The displacement
S=(v_0+v)/2 t=(0+460)/2×2.75×〖10〗^(-3)=0.6325 m=63.3 cm